OpenStudy (anonymous):

Given the point (4, 5) and the slope of 6, find y when x = 24.

jimthompson5910 (jim_thompson5910):

plug in m = 6, x = 4 and y = 5 then solve for b y = mx+b y = 6x + b 5 = 6*4 + b ... ... ... b = ??

OpenStudy (anonymous):

@jim_thompson5910 confused.. my answer choices are 118 144 125 149

jimthompson5910 (jim_thompson5910):

don't worry about the final answer, let's focus on solving for b first 5 = 6*4 + b 5 = 24 + b b = ??

OpenStudy (anonymous):

@jim_thompson5910 -19

jimthompson5910 (jim_thompson5910):

so y = mx+b turns into y = 6x - 19

jimthompson5910 (jim_thompson5910):

now plug in x = 24 y = 6x - 19 y = 6*24 - 19 y = ???

OpenStudy (anonymous):

@jim_thompson5910 125 thank you!